One eigth of the square root of 7 less than a number is 2. What is the number?
Answer by checkley77(12844)

x-1/8(sqrt7)=2
x=2+1/8(sqrt7)
x=2+1/8*2.64575
x=2+.3307
x=2.3307 ans.